3x-4y=7
2x-y=3
then
3x-4y=7
-8x+4y=-12
adding
-5x=-5 then x=1 then 3-4y=7 then 4y=-4 then y=-1
Solution: x=1,y=-1

Answer by oscargut(2103)

3x-4y=7
2x-y=3
then
3x-4y=7
-8x+4y=-12
adding
-5x=-5 then x=1 then 3-4y=7 then 4y=-4 then y=-1
Solution: x=1,y=-1